OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 2x − y = 4. x-intercept: (−4, 0), y-intercept: (0, 2) x-intercept: (−4, 0), y-intercept: (2, 0) x-intercept: (2, 0), y-intercept: (0, −4) x-intercept: (0, 2), y-intercept: (0, −4)

OpenStudy (shamil98):

2x - y = 4 Write this in slope-intercept form, and you will have your answer.

OpenStudy (shamil98):

y = mx + b

OpenStudy (anonymous):

2x-4=y

HanAkoSolo (jamierox4ev3r):

2x-y=4 -2x -2x ----- -y=-2x+4 y=2x-4

HanAkoSolo (jamierox4ev3r):

if the slope intercept form of the equation is y=2x-4, what is the slope and the y-intercept @shahbaaz786 ?

OpenStudy (anonymous):

slope=2 and then 4

OpenStudy (anonymous):

is that right ?

HanAkoSolo (jamierox4ev3r):

yes.... 4 is the y-intercept ;)

OpenStudy (anonymous):

so thats the answer ?

HanAkoSolo (jamierox4ev3r):

based from what i told you, you have to figure that out :) I don't give answers, I just tell you how to do it

OpenStudy (anonymous):

so how do i solve it ?

OpenStudy (shamil98):

b from the y =mx + b is the y intercept in this problem it is -4. >_>

OpenStudy (anonymous):

but theres 2 with -4

HanAkoSolo (jamierox4ev3r):

@shamil98 I will let you take it from here..... lol

OpenStudy (shamil98):

nooo you do ;c

OpenStudy (shamil98):

I was gonna say that haha

HanAkoSolo (jamierox4ev3r):

well i said it to... finder's keepers XD

OpenStudy (shamil98):

Graph it and find the x intercept using slope ._.

OpenStudy (anonymous):

cant i just plug in

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Machida

OpenStudy (anonymous):

@uri